# WixUp: A General Data Augmentation Framework for Wireless Perception in Tracking of Humans

### v1

@@ 5.3.1 Unsupervised domain adaptation across users
In the context of user domain adaptation, the steps of UDA via self-training are the same as the above 4.3. In step 1, the model undergoes training using labeled data sourced from the training users, constituting the source domain. Subsequently, in step 2, the device is sold to a new user, who starts by inputting some new data designated for self-training. Moving to step 3, the trained model performs inference on the unlabeled data from the new user, generating output predictions that are retained as pseudo-labels. Step 4 involves pairing one sample from the source domain data with one from the new user data, and mixing them via WixUp to form semi-labeled training data. Finally, in step 5, the newly created training data is utilized to fine-tune the trained model. Consequently, the model’s performance on the future new user data is expected to exhibit improvement. In other words, the training and labeling are confined to the efforts of the sensor developer; when a user buys a new device, they can effortlessly improve the device performance without labels of their data, since the device might not equipped with ground truth cameras.
@@
5.3.2 Unsupervised domain adaptation across environment
@@
To validate this, we use the MMFi dataset because it has clearly labeled four scenes in data collection. Note that new scenes also mean new users in this dataset. Therefore, we train and evaluate UDA across environments in a leave-one-scene-out manner for the four scenes. The right in Fig. 4 depicts the error of MLE in cm for keypoint pose estimation. All four scenes yield a large margin of performance improvement with (w/) WixUp than that without (w/o) WixUp. The average improvements are 17.45%, ranging from 15.04% to 21.73%.
@@
5.4 Ablation study of WixUp
@@
Figure 5: Expanding the augment data size by tweaking WixUp brings further drops in errors.
@@
Table 5: Ablation study on incremental versions of WixUp proves the effectiveness of each component.
5.4.1 Augmentation size
First, we investigate the impact of mixing distance, which refers to the number of interval frames between the two samples to be mixed. For example, distance=2 means each pair of mixing data is sampled from one real data along with its neighbor located two frames ahead. Usually, stacking pairs from varying distances results in a larger augmentation size and thus greater diversity in data distribution. Although augmenting more data typically improves accuracy, there may be a turning point or plateau point where improvements slow down. As shown in Fig. 5, we increase the distance from one to ten, in the context of key point estimation in MiliPoint with the error as MLE in cm. Excessive augmentation makes the benefits slow down but, impressively, it is still decreasing; The slow down is possibly due to the decreased accuracy in mixing distant pairs. It could also be because the benefits of enriching data distribution reach a limit, ceasing to other primary bottlenecks such as factors in learning algorithms. To clarify, the above experiments only augment by a distance of one by default, in order to facilitate the experiments; thereby, we could expect more drops in their error when excessively fine-tuning a single job.
5.4.2 Effectiveness of algorithmic components
@@
6 Discussion
In this paper, our focus is on human tracking tasks within wireless perception, because this popular field provides enough open-source large datasets and methods ready for us to conduct comprehensive DA research. However, it’s important to acknowledge that wireless perception encompasses a broader spectrum of applications beyond human tracking, for instance, healthcare monitoring, smart agriculture, and supply chain management. We hope to extend WixUp to these applications once they grow up with more resources in the future.
Moreover, WiFi is also a promising modality for wireless applications, many of which adhere to similar FMCW-based modulation and processing pipelines. However, the costly hardware and excessive installation efforts make this area have less open-source data in public. For instance, [24, 62, 61] are notable works on this topic but only published the code. We hope WixUp can embark on open research in this modality as well.

### v2

@@ 5.3.1. Unsupervised domain adaptation across users
In the context of user domain adaptation, the steps of UDA via self-training are the same as the above 4.3. 1) The model undergoes training using labeled data sourced from the training users, constituting the source domain. 2) The device is sold to a new user, who starts by inputting some new data designated for self-training. 3) The trained model performs inference on the unlabeled data from the new user, generating output predictions that are retained as pseudo-labels. 4) Pair one sample from the source domain data with one from the new user data, and mix them via WixUp to form semi-labeled training data. 5) The newly created training data is utilized to fine-tune the trained model. Consequently, the model’s performance on the future new user data is expected to exhibit improvement. In other words, the training and labeling are confined to the efforts of the sensor developer; when a user buys a new device, they can effortlessly improve the device performance without labels of their data, since the device might not equipped with ground truth cameras.
@@
5.3.2. Unsupervised domain adaptation across environment
@@
To validate this, we use the MMFi dataset, which has four clearly labeled scenes in data collection. Note that new scenes also imply new users. Therefore, we train and evaluate UDA across scenes in a leave-one-scene-out manner. The right in Fig. 4 depicts the error of MLE in cm for keypoint pose estimation. All four scenes yield a large margin of performance improvement with (w/) WixUp than that without (w/o) WixUp. The average improvements are 17.45%, ranging from 15.04% to 21.73%.
@@
5.3.3. Comparison with generative-model-based DA
Table 5 summarizes a comprehensive comparison with generative-model-based DA by reporting their experiment results and applicable tasks, showing absolute improvements in percentages and relative improvements (in parentheses). NA means the model is not applicable to that task; dash - means theoretically applicable but not reported in their experiments. To make the comparison relatively fair despite differences in datasets or tasks, we report a range of all results in this scenario.
All these works test on their own collected dataset, focusing on RF signals. Each of them is confined to certain tasks and dataset, making them invalid for a comprehensive comparison to our benchmark results as baselines. Most methods could facilitate cross-domain adaptation. However, RF-Diffusion (chi2024rfdiff, ) was only tested on action classification tasks, as it relies on conditional generation, which may face challenges with longer condition vectors, such as sequences of 99 coordinates, which have yet to be validated. RF-Gen (chen2023rfgenesis, ) demonstrated strong improvements in domain adaptation but cannot be applied to identification tasks, as its text prompt cannot depict this complex feature. NeRF2 (zhao2023nerf2, ), focused on static scene reconstruction, has no experiments in dynamic or cross-domain scenarios where its method would significantly degrade. In general, these generative methods are limited to certain tasks or data, whose experiments are on their own collected data and compare with one no-synthesis baseline, or up to three other generative models. In contrast, WixUp demonstrates strong generalization to various tasks, along with validation across a variety of datasets, model architectures, modalities, and baseline comparisons as above; moreover, our domain adaptation can even be achieved in an unsupervised manner through self-training, which produces higher-quality semi-labeled synthetic samples.
Action (body) Identification
Domain adaptation
(Action/Keypoint)
NeRF2 (zhao2023nerf2, ) NA NA NA
RF-Gen (chen2023rfgenesis, ) - NA (6.04$\sim$56.42%)
RF-Diffusion (chi2024rfdiff, ) 1.8$\sim$8.7%(1.98$\sim$11.01%) - (5.34$\sim$16.20%)
WixUp 11%(84.25%) 2.0% (5.29%) (1.26$\sim$21.73%)
Table 5. Comprehensive comparison with generative DA. [NA = not applicable; - = not reported]
5.4. Ablation study of WixUp
@@
Figure 5. Expanding the augment data size by tweaking WixUp brings further drops in errors.
@@
Table 6. Ablation study on incremental versions of WixUp proves the effectiveness of each component.
5.4.1. Augmentation size
First, we investigate the impact of mixing distance, which refers to the number of interval frames between the two samples to be mixed. For example, distance=2 means each pair of mixing data is sampled from one real data along with its neighbor located two frames ahead. When increasing the maximum allowable distance, data augmentation scale also increases. Scale refers to the factor by which the augmentation data size exceeds the original data size. For instance, scale=6 means we mix pairs whose distance=1,2,3,4,5,or 6; the mixing at each distance yields new data equivalent in size to the original data (ignoring the trivial number of unreachable pairs at the tail, around only distance-1 pairs), thus scale=6 implies that the augmentation data is 6 times the size of the original. Usually, stacking pairs from varying distances results in a larger augmentation size and thus greater diversity in data distribution. Although augmenting more data typically improves accuracy, there may be a turning point or plateau point where improvements slow down. As shown in Fig. 5, we increase the distance from one to ten,  that is equivalent to a maximum distance of around 0.3s. It is in the context of key point estimation in MiliPoint with the error as MLE in cm. Excessive augmentation makes the benefits slow down when scale¿6 but, impressively, it is still decreasing; The slow down is possibly due to the decreased accuracy in mixing distant pairs. It could also be because the benefits of enriching data distribution reach a limit, ceasing to other primary bottlenecks such as learning algorithms’ limited scaling law. To clarify, the above experiments only augment by a distance of one by default, in order to facilitate the experiments; thereby, we could expect more drops in their error when excessively fine-tuning a single job. Besides, since we used a 20% subset of real data for the ablation experiments, we demonstrate that increasing real data (up to three additional scales, excluding the test data) reduces error similarly. This shows our augmentation data has equivalent realness whereas augmentation can be scaled up to three times or more, unlimited to the amount of real data available. In general, to select an optimal augmentation, developers should consider the dataset frame rate, downstream model and task, as well as the computation resource constraints.
5.4.2. Effectiveness of algorithmic components
@@
6. Discussion
In this paper, we focus on human tracking tasks within wireless perception, since its popularity brings ample open-source large datasets and methods, ready for us to conduct comprehensive DA research. However, it’s important to acknowledge that wireless perception encompasses a broader spectrum of applications beyond human tracking, for instance, healthcare monitoring, smart agriculture, and supply chain management. We hope to extend WixUp to these applications once they mature with more resources in the future.
Moreover, WiFi is also a promising modality for wireless applications, many of which adhere to similar FMCW-based modulation and processing pipelines. However, the costly hardware and excessive installation efforts make this area have less open-source data in public. For instance, (jiang2020towards, ; zhao2018rf, ; zhao2018through, ) are notable works on this topic but only published the code. We hope WixUp can embark on open research in this modality as well.
@@ 6. Discussion
As mentioned in Sec. 4.1, our approach involves mixing the joint-wise positions in the midway. This could produce infeasible joint positions if the frame rate is low or the original samples are not close in the time sequences. While these issues usually do not occur in supervised learning use cases, they could be pronounced during unsupervised domain adaptation. Because, in our experiments of UDA via self-training, mixing involves samples from distinct time sequences, such as two samples from distinct users or environments. This might produce infeasible joint positions, such as shortened or intersected arms. Further ablation studies might help understand this effect comprehensively. Nevertheless, WixUp still achieved significant improvements in UDA. We think this could be attributed to the fact that the theoretical foundation of convex combination  (zhang2017mixup, ) can work well without requiring stringent preconditions. Specifically, in convex combination, the mixing data augmentation encourages the model to behave linearly in between training examples. This linear behavior reduces the amount of undesirable oscillations in vicinal data distribution. Vicinal distribution can be noisy, for instance, traditional approaches even simply generate augmentation data from Gaussian noise. For reference,  (zhang2017mixup, ) mixes images for classification, such as mixing images of dogs and cats pixel-wise, producing a new noisy image of an infeasible creature; but they also have shown the effectiveness of mixing-based data augmentation in improving data distribution and model performance. Furthermore, in future work, for both supervised learning and the UDA cases, we can adopt an add-on filter using forward kinematic constraints to alleviate this potential issue.